# Open Heavy flavour mesons in hot asymmetric strange hadronic matter – a QCD sum rule approach

### v1

@@ II QCD sum rule approach
where, $f=\Pi_{OPE}^{e}(M^{2})$ and $g=\Pi_{OPE}^{o}(M^{2})$, the even and odd parts of the spectral function, are given by equations (5) and (7), and, $f^{\prime}$ and $g^{\prime}$ denote the derivatives of $f$ and $g$ with respect to $1/{M^{2}}$. As can be observed from eqn (10), the mass difference between the particles ($D^{0}$, $D^{+}$ and $D_{s}^{+}$) and their antiparticles ($\bar{D^{0}}$, ${D^{-}}$ and $D_{s}^{-}$) for the charm sector, and between the particles ($B^{-}$, $\bar{B^{0}}$ and ${\bar{B_{s}}}^{0}$) and their antiparticles ($B^{+}$, $B^{0}$ and $B_{s}^{0}$) for the bottom sector, in the hadronic medium arise from the values of $\langle{q_{i}}^{\dagger}{q_{i}}\rangle$, $\langle{q_{i}}^{\dagger}iD_{0}^{2}{q_{i}}\rangle$ and $\langle{q_{i}}^{\dagger}g_{s}\sigma.G{q_{i}}\rangle$ of the odd part of the spectral function ($g$), hence due to the expectation values of the QCD operators $\langle{q_{i}}^{\dagger}{q_{i}}\rangle$, $\langle{q_{i}}^{\dagger}D_{0}^{2}{q_{i}}\rangle$ and $\langle{q_{i}}^{\dagger}g_{s}\sigma\cdot Gq_{i}\rangle$ in $P^{o}(M^{2})$ given by equation (8). The in-medium QCD condensates are calculated within a chiral SU(3) model, as described in the following section.
@@ III Chiral SU(3) model
In this section, we briefly describe the chiral SU(3) model paper3 ; hartree ; kristof1 ; AMAKSPM24 which is used to obtain the in-medium QCD condenstaes in the hot asymmetric strange hadronic matter. The expectation values of these QCD operators are central to the calculation of the in-medium masses of the open charm and open bottom mesons within the framework of the QCD sum rule approach, as described in the previous section. The chiral SU(3) model paper3 ; hartree ; kristof1 is based on a non-linear realization of chiral symmetry Weinberg ; coleman ; Bardeen and incorporates the broken scale invariance of QCD through introduction of a scalar dilaton field, $\chi$ sche1 ; heide1 . The Lagrangian density of chiral SU(3) model is written as paper3
@@
In Eq.(11), ${\cal L}_{kin}$ is the kinetic energy term, ${\cal L}_{BW}$ is the baryon-meson interaction term, where, $W=X,Y,V,A,u$ correspond to the interactions of the baryons with the scalar, pseudoscalar singlet, vector meson, axial-vector meson and pseudoscalar octet mesons respectively. The baryons-spin-0 meson interaction term generates the baryon masses. ${\cal L}_{vec}$ describes the dynamical mass generation of the vector mesons via couplings to the scalar mesons and contain additionally quartic self-interactions of the vector fields. ${\cal L}_{0}$ contains the meson-meson interaction terms, ${\cal L}_{\chi}$ is the QCD scale breaking term in the Lagrangian density, including a logarthimic potential of the scalar dilaton field, $\chi$, and, ${\cal L}_{SB}$ describes the explicit chiral symmetry breaking. In the present study of isospin asymmetric strange hadronic matter at finite temperature, the mean field approximation is used, in which the meson fields are replaced by their expectation values. The meson fields which have non-zero expectation values are the scalar ($\sigma,\zeta,\delta)$ and the vector fields, $\omega^{\mu}\rightarrow\omega\delta^{\mu 0}$, $\rho^{\mu a}\rightarrow\rho\delta^{\mu 0}\delta_{a3}$ and, $\phi^{\mu}\rightarrow\phi\delta^{\mu 0}$, and, the expectation values of the other mesons are zero. We also use the approximations $\bar{\psi}_{i}\psi_{j}\rightarrow\delta_{ij}\langle\bar{\psi}_{i}\psi_{i}\rangle\equiv\delta_{ij}\rho_{i}^{s}$ and $\bar{\psi}_{i}\gamma^{\mu}\psi_{j}\rightarrow\delta_{ij}\delta^{\mu 0}\langle\bar{\psi}_{i}\gamma^{0}\psi_{i}\rangle\equiv\delta_{ij}\delta^{\mu 0}\rho_{i}$, where, $\rho_{i}^{s}$ and $\rho_{i}$ are the scalar and number density of baryon of species, $i$. In the above approximation, the terms which contribute to the baryon-meson interaction, ${\cal L}_{BW}$ of equation (11) arise from the interactions due to the scalar ($S$) and the vector ($V$) mesons, and, is given as
@@
Figure 1: Masses of $D^{\pm}$ (in panels (a), (c) and (e)) and $D^{0}(\bar{D^{0}})$ (in panels (b), (d) and (f)) mesons are plotted as functions of the Borel mass M for T=50, 100 and 150 MeV in nuclear matter ($f_{s}$=0) for baryon density $\rho_{B}=2\rho_{0}$ and $\eta=0.5$, and compared to the case of $\eta$=0 (shown as dotted lines). The Borel curve for $\rho_{B}=0$ is shown as the dot-dashed line.
Figure 2: Same as Fig. 1, with $f_{s}$=0.3.
Figure 3: Mass shifts of $D^{\pm}$ (in panels (a), (c) and (e)) and $D^{0}(\bar{D^{0}})$ (in panels (b), (d) and (f)) mesons are plotted as functions of the baryon density in units of nuclear saturation density for T=50, 100 and 150 MeV in nuclear matter ($f_{s}$=0) for $\eta=0.5$, and compared to the case of $\eta$=0 (shown as dotted lines).
Figure 4: Same as fig. 3, with $f_{s}$=0.3.
@@
Figure 5: Masses of $D_{s}^{\pm}$ mesons are plotted as functions of the Borel mass M for T=50, 100 and 150 MeV in panels (a), (c) and (e) and in panels (b), (d) and (f), in nuclear ($f_{s}$=0) and hyperonic (with $f_{s}$=0.3) matter for baryon density $\rho_{B}=2\rho_{0}$ and $\eta=0.5$, and compared to the case of $\eta$=0 (shown as dotted lines). The Borel curve for $\rho_{B}=0$ is shown as the dot-dashed line.
Figure 6: Mass shifts of $D_{s}^{\pm}$ mesons are plotted as functions of the baryon density in units of nuclear saturation density for $\eta=0.5$ (in panels (a), (c) and (e)) for nuclear matter ($f_{s}$=0) and in panels (b), (d) and (f) for strange hadronic matter with $f_{s}$=0.3) for T=50, 100 and 150 MeV and compared to the case of $\eta$=0 (shown as dotted lines).
Figure 7: Masses of $B^{+}({B^{-}})$ (in panels (a), (c) and (e)) and $B^{0}(\bar{B}^{0})$ (in panels (b), (d) and (f)) mesons are plotted as functions of the Borel mass M for T=50, 100 and 150 MeV in nuclear matter ($f_{s}$=0) for baryon density $\rho_{B}=2\rho_{0}$ and $\eta=0.5$, and compared to the case of $\eta$=0 (shown as dotted lines). The Borel curve for $\rho_{B}=0$ is shown as the dot-dashed line.
Figure 8: Same as fig. 7, with $f_{s}$=0.3.
@@
Figure 9: Mass shifts of $B^{+}({B^{-}})$ (in panels (a), (c) and (e)) and $B^{0}(\bar{B}^{0})$ (in panels (b), (d) and (f)) mesons are plotted as functions of the baryon density in units of nuclear saturation density for T=50, 100 and 150 MeV in nuclear matter ($f_{s}$=0) for $\eta=0.5$, and compared to the case of $\eta$=0 (shown as dotted lines).
Figure 10: Same as fig. 9, with $f_{s}$=0.3.
Figure 11: Masses of $B_{s}^{0}$ and ${\bar{B}}_{s}^{0}$ mesons are plotted as functions of the Borel mass M for T=50, 100 and 150 MeV for nuclear matter ($f_{s}$=0) in panels (a), (c) and (e) and for hyperonic (with $f_{s}$=0.3) in panels (b), (d) and (f), for baryon density $\rho_{B}=2\rho_{0}$ and $\eta=0.5$, and compared to the case of $\eta$=0 (shown as dotted lines). The Borel curve for $\rho_{B}=0$ is shown as the dot-dashed line.
@@
Figure 12: Mass shifts of $B_{s}^{0}$ and ${\bar{B}}_{s}^{0}$ mesons are plotted as functions of the baryon density in units of nuclear saturation density for $\eta=0.5$ (in panels (a), (c) and (e)) for nuclear matter ($f_{s}$=0) and in panels (b), (d) and (f) for strange hadronic matter with $f_{s}$=0.3) for T=50, 100 and 150 MeV and compared to the case of $\eta$=0 (shown as dotted lines).
@@
The parameter $d$ in equation (20) originates from the second logarithmic term of equation (15). The parameter $d$, along with the other parameters corresponding to the meson-meson interaction Lagrangian density, ${\cal L}_{0}$ of the general Lagrangian given by (11), are fitted so as to ensure extrema in the vacuum for the $\sigma$, $\zeta$ and $\chi$ field equations, to reproduce the vacuum masses of the $\eta$ and $\eta^{\prime}$ mesons, the mass of the $\sigma$ meson around 500 MeV, and, pressure, p($\rho_{0}$)=0, with $\rho_{0}$ as the nuclear matter saturation density paper3 ; amarindamprc . The scalar fields (the non-strange isoscalar field, $\sigma$, the strange isoscalar field, $\zeta$ and the isovector field $\delta$), the dilaton field, $\chi$ and the vector fields (non-strange isoscalar $\omega$, strange isoscalar $\phi$ and non-strange isovector $\rho$) are obtained by minimizing the thermodynamic potential $\Omega$ for given values of temperature, baryonic density, $\rho_{B}$, the isospin asymmetry parameter, $\eta=-\frac{\Sigma_{i}I_{3i}\rho_{i}}{\rho_{B}}$, the strangeness fraction, $f_{s}=\frac{\Sigma_{i}|s_{i}|\rho_{i}}{\rho_{B}}$, where, $I_{3i}$ and $|s_{i}|$ denote the 3rd component of isospin quantum number and number of strange quarks in $i^{th}$ baryon, respectively.
The effects of the scalar and twist-2 gluon condensates, as obtained from the medium modification of the dilaton field, $\chi$, within the chiral SU(3) model, on the masses of the vector and pseudoscalar charmonium ground states ($J/\psi$ and $\eta_{c}$) were studied in Ref. amarvjpsi_qsr and on the masses of the heavy quarkonium (charmonium and bottomonium) states in Ref. QCDSR_Heavy_Quarkonia , using QCD sum rule approach. In the present work, the effects of the medium modifications of the QCD operators, upto dimension 5, on the masses of the open charm and open bottom mesons are investigated within the QCD sum rule approach. The non-strange ($\langle\bar{u}u\rangle$, $\langle\bar{d}d\rangle$) and strange ($\langle\bar{s}s\rangle$) quark condensates are obtained from the scalar fields using equation (14) and the scalar and twist-2 gluon condensates are calculated from the dilaton field, $\chi$, using equations (20 and (21) within the chiral SU(3) model. The dimension 5 mixed quark-gluon condensate $\langle\bar{q_{i}}g_{s}\sigma.G{q_{i}}\rangle$ in the operator product expansion is obtained from the in-medium quark condensate $\langle\bar{q_{i}}{q_{i}}\rangle$, with $i=1,2,3\equiv u,d,s$, as computed within the chiral SU(3) model, by using the relation $\langle\bar{q_{i}}g_{s}\sigma.G{q_{i}}\rangle=2\lambda_{q}^{2}\langle\bar{q_{i}}{q_{i}}\rangle$. The estimates for $\lambda_{q}^{2}$ Cohen_Prog_Part_Nucl_Phys in vacuum using standard QCD sum rule approach is 0.4 GeV2 Belyaev , the value estimated using lattice QCD is 0.55 GeV2 Kremer , and, $\lambda_{q}^{2}\sim$ 1.2 GeV2, in a QCD sum rule analysis of the pion wave function using non-local condensates Mikhailov as well as within an instanton liquid model Shuryak . Using the leading order approximations for the parton distribution functions Glueck_Reya_Vogt , the values of the higher order operators (or combinations of these operators) used in the present investigation are as follows Jin_Cohen_Furnstahl_Griegel_PRC47_2882_1983 . The estimates for $q_{i}=u,d$ are
@@
for the case of symmetric nuclear matter. For the strange quark ($q_{i}=s$), the estimate used in the present investigation for the above operator is taken to be $\langle s^{\dagger}iD_{0}s\rangle=0.06\;{\rm GeV}\rho_{B}$ Cohen_Prog_Part_Nucl_Phys . For the consideration of the strange-heavy flavor mesons, the estimates of the combinations of the operators given by equations (25) and (26) (for $q_{i}$=s) are given as open_heavy_flavour_qsr_2
@@
$y=\frac{\langle\bar{s}s\rangle_{med}}{\langle\bar{q}q\rangle_{med}},$ (29)
with $\langle\bar{s}s\rangle_{med}=\langle\bar{s}s\rangle-\langle\bar{s}s\rangle_{vac}$, and, ${\langle\bar{q}q\rangle}_{med}=({\langle\bar{u}u\rangle}_{med}+{\langle\bar{d}d\rangle}_{med})/2$ is the average of the quark-antiquark condensates for $u$ and $d$ quarks after subtracting the vacuum values. For the isospin asymmetric nuclear (hyperonic) matter as considered in the present investigation, we generalize the above relations as follows. For the non-strange quarks, $q_{i}=u,d$, for $i=1,2$,
$-\langle{\bar{q}_{i}}D_{0}^{2}q\rangle+\frac{1}{8}\langle\bar{q_{i}}g_{s}\sigma.G{q_{i}}\rangle=0.3\;{\rm GeV}^{2}\;\rho_{B}\cdot\frac{\langle\bar{q_{i}}q_{i}\rangle_{med}}{\langle\bar{q}q\rangle_{med}},$ (30)
$-\langle q_{i}^{\dagger}D_{0}^{2}q_{i}\rangle+\frac{1}{12}\langle{q_{i}}^{\dagger}g_{s}\sigma.G{q_{i}}\rangle=0.031\;{\rm GeV}^{2}\;\rho_{B}\cdot\frac{\langle\bar{q_{i}}q_{i}\rangle_{med}}{\langle\bar{q}q\rangle_{med}},$ (31)
@@
$-\langle{\bar{s}}D_{0}^{2}s\rangle+\frac{1}{8}\langle\bar{s}g_{s}\sigma.G{s}\rangle=0.3\;{\rm GeV}^{2}\;\rho_{B}\cdot y,$ (32)
$-\langle s^{\dagger}D_{0}^{2}s\rangle+\frac{1}{12}\langle{s}^{\dagger}g_{s}\sigma.G{s}\rangle=0.031\;{\rm GeV}^{2}\;\rho_{B}\cdot y,$ (33)
with $y$, the strangeness content of the nucleon defined by equation (29). The estimate for the operator $\langle q_{i}^{\dagger}g_{s}\sigma.Gq_{i}\rangle=0.33\;{\rm{GeV}}^{2}\rho_{B}$, for $q_{i}=u\;\;{\rm{or}}\;\;d$ for the symmetric nuclear matter open_heavy_flavour_qsr_2 is generalized to the isospin asymmetric case in the present investigation as
$\langle q_{i}^{\dagger}g_{s}\sigma.Gq_{i}\rangle=0.33\;{\rm{GeV}}^{2}\rho_{B}\cdot\frac{\langle\bar{q_{i}}q_{i}\rangle_{med}}{\langle\bar{q}q\rangle_{med}},$ (34)

### v2

@@ II QCD sum rule approach
where, $f=\frac{1}{\pi}\int_{m_{Q}^{2}}^{{\tilde{s}}_{0}}{\rm d}{\tilde{s}}{\rm Im}\;\Pi_{OPE}^{per}({\tilde{s}},\vec{q}=0)e^{-{\tilde{s}}/M^{2}}+e^{-{m_{Q}^{2}}/{M^{2}}}P^{e}(M^{2})$, with $P^{e}(M^{2})$ given by equation (7), $g=e^{-{m_{Q}^{2}}/{M^{2}}}P^{o}(M^{2})$, the odd part of the spectral function as given by equation (8), and, $f^{\prime}$ and $g^{\prime}$ denote the derivatives of $f$ and $g$ with respect to $1/{M^{2}}$. As can be observed from eqn (10), the mass difference between the particles ($D^{0}$, $D^{+}$ and $D_{s}^{+}$) and their antiparticles ($\bar{D^{0}}$, ${D^{-}}$ and $D_{s}^{-}$) for the charm sector, and between the particles ($B^{-}$, $\bar{B^{0}}$ and ${\bar{B_{s}}}^{0}$) and their antiparticles ($B^{+}$, $B^{0}$ and $B_{s}^{0}$) for the bottom sector, in the hadronic medium arise from the values of $\langle{q_{i}}^{\dagger}{q_{i}}\rangle$, $\langle{q_{i}}^{\dagger}iD_{0}^{2}{q_{i}}\rangle$ and $\langle{q_{i}}^{\dagger}g_{s}\sigma.G{q_{i}}\rangle$ of the odd part of the spectral function ($g$), hence due to the expectation values of the QCD operators $\langle{q_{i}}^{\dagger}{q_{i}}\rangle$, $\langle{q_{i}}^{\dagger}D_{0}^{2}{q_{i}}\rangle$ and $\langle{q_{i}}^{\dagger}g_{s}\sigma\cdot Gq_{i}\rangle$ in $P^{o}(M^{2})$, given by equation (9). The in-medium QCD condensates are calculated within a chiral SU(3) model, as described in the following section.
@@ III Chiral SU(3) model
In this section, we briefly describe the chiral SU(3) model paper3 ; hartree ; kristof1 ; AMAKSPM24 which is used to obtain the in-medium QCD condenstaes in the hot asymmetric strange hadronic matter. The expectation values of these QCD operators are central to the calculation of the in-medium masses of the open charm and open bottom mesons within the framework of the QCD sum rule approach, as described in the previous section. The chiral SU(3) model paper3 ; hartree ; kristof1 is based on a non-linear realization of chiral symmetry Weinberg ; coleman ; Bardeen and incorporates the broken scale invariance of QCD through the introduction of a scalar dilaton field, $\chi$ sche1 ; heide1 . The Lagrangian density of chiral SU(3) model is written as paper3
@@
In Eq.(11), ${\cal L}_{kin}$ is the kinetic energy term, ${\cal L}_{BW}$ is the baryon-meson interaction term, where, $W=X,Y,V,A,u$ correspond to the interactions of the baryons with the scalar, pseudoscalar singlet, vector meson, axial-vector meson and pseudoscalar octet mesons respectively. The baryons-spin-0 meson interaction term generates the baryon masses. ${\cal L}_{vec}$ describes the dynamical mass generation of the vector mesons via couplings to the scalar mesons and contain additionally quartic self-interactions of the vector fields. ${\cal L}_{0}$ contains the meson-meson interaction terms, ${\cal L}_{\chi}$ is the QCD scale breaking term in the Lagrangian density, including a logarthimic potential of the scalar dilaton field, $\chi$, and, ${\cal L}_{SB}$ describes the explicit chiral symmetry breaking. In the present study of isospin asymmetric strange hadronic matter at finite temperature, the mean field approximation is used, in which the meson fields are replaced by their expectation values. The meson fields which have non-zero expectation values are the scalar ($\sigma,\zeta,\delta)$ and the vector fields, $\omega^{\mu}\rightarrow\omega\delta^{\mu 0}$, $\rho^{\mu a}\rightarrow\rho\delta^{\mu 0}\delta^{a3}$ and, $\phi^{\mu}\rightarrow\phi\delta^{\mu 0}$, and, the expectation values of the other mesons are zero. We also use the approximations $\bar{\psi}_{i}\psi_{j}\rightarrow\delta_{ij}\langle\bar{\psi}_{i}\psi_{i}\rangle\equiv\delta_{ij}\rho_{i}^{s}$ and $\bar{\psi}_{i}\gamma^{\mu}\psi_{j}\rightarrow\delta_{ij}\delta^{\mu 0}\langle\bar{\psi}_{i}\gamma^{0}\psi_{i}\rangle\equiv\delta_{ij}\delta^{\mu 0}\rho_{i}$, where, $\rho_{i}^{s}$ and $\rho_{i}$ are the scalar and number density of baryon of species, $i$. In the above approximation, the terms which contribute to the baryon-meson interaction, ${\cal L}_{BW}$ of equation (11) arise from the interactions due to the scalar ($S$) and the vector ($V$) mesons, and, is given as
@@
Figure 1: Masses of $D^{+}(D^{-})$ (in panels (a), (c) and (e)) and $D^{0}(\bar{D^{0}})$ (in panels (b), (d) and (f)) mesons are plotted as functions of the Borel mass M for T=50, 100 and 150 MeV in nuclear matter ($f_{s}$=0) for baryon density $\rho_{B}=2\rho_{0}$ and $\eta=0.5$, as the solid (dashed) lines. These are compared to the case of $\eta$=0 (shown as dotted lines). The Borel curve for $\rho_{B}=0$ is also shown in the figure as the dot-dashed line.
@@
Figure 2: Masses of $D^{+}(D^{-})$ (in panels (a), (c) and (e)) and $D^{0}(\bar{D^{0}})$ (in panels (b), (d) and (f)) mesons are plotted as functions of the Borel mass M for T=50, 100 and 150 MeV in hyperonic matter (with $f_{s}$=0.3) for baryon density $\rho_{B}=2\rho_{0}$ and $\eta=0.5$, as the solid (dashed) lines. These are compared to the case of $\eta$=0 (shown as dotted lines). The Borel curve for $\rho_{B}=0$ is also shown in the figure as the dot-dashed line.
@@
Figure 3: Mass shifts of $D^{+}(D^{-})$ (in panels (a), (c) and (e)) and $D^{0}(\bar{D^{0}})$ (in panels (b), (d) and (f)) mesons are plotted as functions of the baryon density in units of nuclear saturation density for T=50, 100 and 150 MeV in nuclear matter ($f_{s}$=0) for $\eta=0.5$, as the solid (dashed) lines, and, are compared to the case of $\eta$=0 (shown as dotted lines).
@@
Figure 4: Mass shifts of $D^{+}(D^{-})$ (in panels (a), (c) and (e)) and $D^{0}(\bar{D^{0}})$ (in panels (b), (d) and (f)) mesons are plotted as functions of the baryon density in units of nuclear saturation density for T=50, 100 and 150 MeV in hyperonic matter (with $f_{s}$=0.3) for $\eta=0.5$, as the solid (dashed) lines, and, are compared to the case of $\eta$=0 (shown as dotted lines).
Figure 5: Masses of $D_{s}^{+}$ and $D_{s}^{-}$ mesons are plotted as functions of the Borel mass M for T=50, 100 and 150 MeV in panels (a), (c) and (e) and in panels (b), (d) and (f), in nuclear ($f_{s}$=0) and hyperonic (with $f_{s}$=0.3) matter for baryon density $\rho_{B}=2\rho_{0}$ and $\eta=0.5$, as the solid (dashed) lines, and, are compared to the case of $\eta$=0 (shown as dotted lines). The Borel curve for $\rho_{B}=0$ is shown as the dot-dashed line.
Figure 6: Mass shifts of $D_{s}^{+}$ and $D_{s}^{-}$ mesons are plotted as functions of the baryon density in units of nuclear saturation density for $\eta=0.5$ (in panels (a), (c) and (e)) for nuclear matter ($f_{s}$=0) and in panels (b), (d) and (f) for strange hadronic matter with $f_{s}$=0.3) for T=50, 100 and 150 MeV as the solid (dashed) lines, and, are compared to the case of $\eta$=0 (shown as dotted lines).
Figure 7: Masses of $B^{-}({B^{+}})$ (in panels (a), (c) and (e)) and $\bar{B}^{0}(B^{0})$ (in panels (b), (d) and (f)) mesons are plotted as functions of the Borel mass M for T=50, 100 and 150 MeV in nuclear matter ($f_{s}$=0) for baryon density $\rho_{B}=2\rho_{0}$ and $\eta=0.5$, as the solid (dashed) lines, and, are compared to the case of $\eta$=0 (shown as dotted lines). The Borel curve for $\rho_{B}=0$ is shown as the dot-dashed line.
Figure 8: Masses of $B^{-}({B^{+}})$ (in panels (a), (c) and (e)) and $\bar{B}^{0}(B^{0})$ (in panels (b), (d) and (f)) mesons are plotted as functions of the Borel mass M for T=50, 100 and 150 MeV in hyperonic matter (with $f_{s}$=0.3) for baryon density $\rho_{B}=2\rho_{0}$ and $\eta=0.5$, as the solid (dashed) lines, and, are compared to the case of $\eta$=0 (shown as dotted lines). The Borel curve for $\rho_{B}=0$ is shown as the dot-dashed line.
Figure 9: Mass shifts of $B^{-}({B^{+}})$ (in panels (a), (c) and (e)) and $\bar{B}^{0}(B^{0})$ (in panels (b), (d) and (f)) mesons are plotted as functions of the baryon density in units of nuclear saturation density for T=50, 100 and 150 MeV in nuclear matter ($f_{s}$=0) for $\eta=0.5$, as the solid (dashed) lines, and, are compared to the case of $\eta$=0 (shown as dotted lines).
Figure 10: Mass shifts of $B^{-}({B^{+}})$ (in panels (a), (c) and (e)) and $\bar{B}^{0}(B^{0})$ (in panels (b), (d) and (f)) mesons are plotted as functions of the baryon density in units of nuclear saturation density for T=50, 100 and 150 MeV in hyperonic matter (with $f_{s}$=0.3) for $\eta=0.5$, as the solid (dashed) lines, and, are compared to the case of $\eta$=0 (shown as dotted lines).
Figure 11: Masses of $\bar{B_{s}^{0}}$ and ${B}_{s}^{0}$ mesons are plotted as function of the Borel mass M for T=50, 100 and 150 MeV for nuclear matter ($f_{s}$=0) in panels (a), (c) and (e) and for hyperonic (with $f_{s}$=0.3) in panels (b), (d) and (f), for baryon density $\rho_{B}=2\rho_{0}$ and $\eta=0.5$, as the solid (dashed) line, and, is compared to the case of $\eta$=0 (shown as dotted line). The Borel curve for $\rho_{B}=0$ is shown as the dot-dashed line.
Figure 12: Mass shifts of $\bar{B_{s}^{0}}$ and $B_{s}^{0}$ mesons are plotted as function of the baryon density in units of nuclear saturation density for $\eta=0.5$ (in panels (a), (c) and (e)) for nuclear matter ($f_{s}$=0) and in panels (b), (d) and (f) for strange hadronic matter with $f_{s}$=0.3) for T=50, 100 and 150 MeV as the solid (dashed) line, and, is compared to the case of $\eta$=0 (shown as dotted line).
Figure 13: Mass shifts of the open charm mesons are plotted as a functions of density for the isospin symmetric ($\eta$=0) nuclear matter ($f_{s}$=0) in panels (a), (c) and (e), and, for the isospin asymmetric (with $\eta$=0.5) nuclear matter ($f_{s}$=0) in panels (b), (d) and (f), for T=50, 100 and 150 MeV respectively. These mass shifts are shown as the short-dash-dotted, dashed, closely spaced dotted, solid, widely spaced dotted, long-dash-dotted for the $D^{-}$, $\bar{D^{0}}$, $D^{0}$, $D^{+}$, $D_{s}^{+}$ and $D_{s}^{-}$ mesons respectively.
Figure 14: Mass shifts of the open charm mesons are plotted as a functions of density for the isospin symmetric ($\eta$=0) hyperonic matter (with $f_{s}$=0.3) in panels (a), (c) and (e), and, for the isospin asymmetric (with $\eta$=0.5) hyperonic matter (with $f_{s}$=0.3) in panels (b), (d) and (f), for T=50, 100 and 150 MeV respectively. These mass shifts are shown as the short-dash-dotted, dashed, closely spaced dotted, solid, widely spaced dotted, long-dash-dotted for the $D^{-}$, $\bar{D^{0}}$, $D^{0}$, $D^{+}$, $D_{s}^{+}$ and $D_{s}^{-}$ mesons respectively.
Figure 15: Mass shifts of the open bottom mesons are plotted as a functions of density for the isospin symmetric ($\eta$=0) nuclear matter ($f_{s}$=0) in panels (a), (c) and (e), and, for the isospin asymmetric (with $\eta$=0.5) nuclear matter ($f_{s}$=0) in panels (b), (d) and (f), for T=50, 100 and 150 MeV respectively. These mass shifts are shown as the short-dash-dotted, dashed, closely spaced dotted, solid, widely spaced dotted, long-dash-dotted for the $B^{0}$, $B^{+}$, $\bar{B^{0}}$, $B^{-}$, $\bar{B_{s}^{0}}$ and $B_{s}^{0}$ mesons respectively.
Figure 16: Mass shifts of the open bottom mesons are plotted as a functions of density for the isospin symmetric ($\eta$=0) hyperonic matter (with $f_{s}$=0.3) in panels (a), (c) and (e), and, for the isospin asymmetric (with $\eta$=0.5) hyperonic matter (with $f_{s}$=0.3) in panels (b), (d) and (f), for T=50, 100 and 150 MeV respectively. These mass shifts are shown as the short-dash-dotted, dashed, closely spaced dotted, solid, widely spaced dotted, long-dash-dotted for the $B^{0}$, $B^{+}$, $\bar{B^{0}}$, $B^{-}$, $\bar{B_{s}^{0}}$ and $B_{s}^{0}$ mesons respectively.
@@
[…]
The effects of the scalar and twist-2 gluon condensates, as obtained from the medium modification of the dilaton field, $\chi$, within the chiral SU(3) model, on the masses of the vector and pseudoscalar charmonium ground states ($J/\psi$ and $\eta_{c}$) were studied in Ref. amarvjpsi_qsr and on the masses of the heavy quarkonium (charmonium and bottomonium) states in Ref. QCDSR_Heavy_Quarkonia , using QCD sum rule approach. In the present work, the effects of the medium modifications of the QCD operators, upto dimension 5, on the masses of the open charm and open bottom mesons are investigated within the QCD sum rule approach. The non-strange ($\langle\bar{u}u\rangle$, $\langle\bar{d}d\rangle$) and strange ($\langle\bar{s}s\rangle$) quark condensates are obtained from the scalar fields using equation (14) and the scalar and twist-2 gluon condensates are calculated from the dilaton field, $\chi$, using equations (20 and (21) within the chiral SU(3) model. The dimension 5 mixed quark-gluon condensate $\langle\bar{q_{i}}g_{s}\sigma.G{q_{i}}\rangle$ in the operator product expansion is obtained from the in-medium quark condensate $\langle\bar{q_{i}}{q_{i}}\rangle$, with $i=1,2,3\equiv u,d,s$, as computed within the chiral SU(3) model, by using the relation
$\langle\bar{q_{i}}g_{s}\sigma.G{q_{i}}\rangle=2\lambda_{q}^{2}\langle\bar{q_{i}}{q_{i}}\rangle.$ (24)
The values of the parameter $\lambda_{q}$ estimated using different methods are, however, observed to be widely different Cohen_Prog_Part_Nucl_Phys ; Belyaev ; lambda_q2_lat_1 ; lambda_q2_lat_2 ; Kremer ; Mikhailov ; Shuryak . Using the leading order approximations for the parton distribution functions Glueck_Reya_Vogt , the values of the higher order operators (or combinations of these operators) used in the present investigation are as follows Jin_Cohen_Furnstahl_Griegel_PRC47_2882_1983 . The estimates for $q_{i}=u,d$ are
@@
for the case of symmetric nuclear matter. For the strange quark ($q_{i}=s$), the estimate for the above operator is taken to be $\langle s^{\dagger}iD_{0}s\rangle=0.06\;{\rm GeV}\rho_{B}$ Cohen_Prog_Part_Nucl_Phys . For the consideration of the strange-heavy flavor mesons, the estimates of the combinations of the operators given by equations (26) and (27) (for $q_{i}$=s) are given as open_heavy_flavour_qsr_2
@@
$y={\langle\bar{s}s\rangle_{med}}/{\langle\bar{q}q\rangle_{med}},$ (30)
with $\langle\bar{s}s\rangle_{med}=\langle\bar{s}s\rangle-\langle\bar{s}s\rangle_{vac}$, and, ${\langle\bar{q}q\rangle}_{med}=({\langle\bar{u}u\rangle}_{med}+{\langle\bar{d}d\rangle}_{med})/2$ is the average of the quark-antiquark condensates for $u$ and $d$ quarks after subtracting the vacuum contributions. For the isospin asymmetric nuclear (hyperonic) matter as considered in the present investigation, we generalize the above relations as follows. For the non-strange quarks, $q_{i}=u,d$, for $i=1,2$,
$-\langle{\bar{q}_{i}}D_{0}^{2}q\rangle+\frac{1}{8}\langle\bar{q_{i}}g_{s}\sigma.G{q_{i}}\rangle=0.3\;{\rm GeV}^{2}\;\bigg{(}\frac{\langle\bar{q_{i}}q_{i}\rangle_{med}}{\langle\bar{q}q\rangle_{med}}\bigg{)}\;\rho_{B},$ (31)
$-\langle q_{i}^{\dagger}D_{0}^{2}q_{i}\rangle+\frac{1}{12}\langle{q_{i}}^{\dagger}g_{s}\sigma.G{q_{i}}\rangle=0.031\;{\rm GeV}^{2}\;\bigg{(}\frac{\langle\bar{q_{i}}q_{i}\rangle_{med}}{\langle\bar{q}q\rangle_{med}}\bigg{)}\;\rho_{B},$ (32)
@@
$-\langle{\bar{s}}D_{0}^{2}s\rangle+\frac{1}{8}\langle\bar{s}g_{s}\sigma.G{s}\rangle=0.3\;{\rm GeV}^{2}\;\rho_{B}\;y,$ (33)
$-\langle s^{\dagger}D_{0}^{2}s\rangle+\frac{1}{12}\langle{s}^{\dagger}g_{s}\sigma.G{s}\rangle=0.031\;{\rm GeV}^{2}\;\rho_{B}\;y,$ (34)
where $y$ is the strangeness content of the nucleon defined by equation (30). The estimate for the operator $\langle q_{i}^{\dagger}g_{s}\sigma.Gq_{i}\rangle=0.33\;{\rm{GeV}}^{2}\rho_{B}$, for $q_{i}=u\;\;{\rm{or}}\;\;d$ for the symmetric nuclear matter open_heavy_flavour_qsr_2 is generalized to the isospin asymmetric case in the present investigation as
$\langle q_{i}^{\dagger}g_{s}\sigma.Gq_{i}\rangle=0.33\;{\rm{GeV}}^{2}\;\bigg{(}\frac{\langle\bar{q_{i}}q_{i}\rangle_{med}}{\langle\bar{q}q\rangle_{med}}\bigg{)}\;\rho_{B},$ (35)